Hello. Could someone please help me with these problems? I don't know what to do.....Thanks:)
1) Determine whether squrt.49/100 is rational or irrational.
2) What is the diagonal length of a TV screen whose dimensions are 80 by 60 cm?
3) How far up a wall will an 11m ladder reach, if the foot of the ladder must be 4m from the base of the wall? Round your answer to the nearest tenth.

4) Find the distance between (–3, –2) and (1, 4)
Thanks!!!!:)

Answer by checkley75(3666)

sqrt49/100=7/10 or .7 or 70% it is very rational.
------------------------------------------------------------
60^2+80^2=x^2
3600+6400=x62
10000=x^2
x=sqrt10000
x=100 answer. length of the diagonal.
------------------------------------------------------------
4^2+x^2=11^2
16+x^2=121
x^2=121-16
x^2=105
x=sqrt105
x=10.25 answer for the height up the wall for this ladder.
-------------------------------------------------------------
the distance between these two points is the diagonal formed by the right triangle with dimentions of:
x=(1+3)=4 & y=4+2)=6 thus usuing the pythagorean theorem a^2+b^2=c^2
4^2+6^2=c^2
16+36=c^2
c^2=52
c=sqrt52
c=7.2111 answer.